# exampleframelinecolor=gray!30, linewidth=0.5pt, backgroundcolor=gray!5, roundcorner=2pt, innertopmargin=8pt, innerbottommargin=8pt, innerleftmargin=8pt, innerrightmargin=8pt, skipabove=skipbelow=

### v2

@@ 3.2 Arithmetic Reasoning
We fine-tune Mistral-7B [24] and Gemma-2 9B [25] on a 50K-sample subset of MetaMathQA [35], and evaluate their performance on GSM8K [36] and MATH [37]. We insert LoRA adapters into all attention projections (query, key, value, and output) as well as both feedforward layers. We report results in Table 2. ABBA achieves superior performance over all other PEFT approaches across both models, and often outperforms full FT.
@@
Method Mistral-7B Gemma-2 9B
# Params GSM8K ($\uparrow$) MATH (($\uparrow$( ↑) # Params GSM8K ($\uparrow$) MATH (($\uparrow$( ↑)
@@ 4.1 Initialization Strategies for ABBA
Initialization of the adapter matrices $B_{1},A_{1}\text{ and }B_{2},A_{2}$ is crucial to ABBA’s performance. A naive LoRA-style initialization, where $B_{1},B_{2}$ are set to zero and $A_{1},A_{2}$ use Kaiming uniform, leads to training failure due to zeroed-out gradients (see Appendix C). To address this, we explore several initialization strategies that combine truncated SVD-based approximations with standard schemes, summarized in Table 3. Inspired by PiSSA [26], one approach is to approximate the base weight $W_{0}$ at initialization. This can be done by initializing one adapter pair using the truncated SVD of $W_{0}$, and the other with scaled constant values (e.g., ones) to prevent gradient explosion. Another strategy initializes both adapter pairs using the top-$r/2$ components from the truncated SVD of $\sqrt{W_{0}}$. A variation of this approach assigns the top-$r/2$ components to one adapter pair and the next-$r/2$ to the other, introducing greater representational diversity and yielding slightly improved results. We also consider a hybrid strategy, where one adapter pair approximates $W_{0}$ via truncated SVD, and the other follows LoRA-style initialization (Kaiming for $A$, zeros for $B$). This configuration performs best and closely resembles the initialization used in HiRA. Our final method adopts this approach: $B_{1},A_{1}$ are initialized using the truncated SVD of $W_{0}$, while $B_{2},A_{2}$ follow LoRA-style initialization.
@@
[…]
$B_{2},A_{2}\leftarrow\text{top-}r/2\text{ from Trunc. SVD}(W_{0})$, $B_{1},A_{1}\leftarrow$ LoRA Init (Zeros, Kaiming) 66.19 18.06
@@ 4.2 Choosing Important Hyperparameters
We established the relationship between ABBA and LoRA scaling in Theorem 2 by comparing their respective factors, $s_{\text{ABBA}}$ and $\alpha_{\text{LoRA}}$. To empirically validate this, we sweep over a range of $\alpha_{\text{LoRA}}$ values and corresponding ABBA scaling factors, ($s_{\text{ABBA}}=\alpha_{\text{LoRA}}^{2}/\sqrt{r_{1}r_{2}})$, for Llama-3.2 3B, in Table 4. Consistent with our theory, ABBA achieves optimal performance within the typical LoRA scaling range of $16-32$. Additional evidence is provided in Table 7 (Appendix E).
Table 4: Performance comparison across different scaling factor values $s_{\text{ABBA}}$ for Llama-3.2 3B.
$\mathbf{s_{\textbf{ABBA}}}$ ($\mathbf{\alpha_{\textbf{LoRA}}}$) Accuracy ($\uparrow$)
@@ 4.3 Placement of ABBA in Transformers
Figure 3 examines the effect of fine-tuning individual transformer components, namely Query, Key, Value, Output, Up, Gate, and Down projections. The results show the following: Query/Key contribute the least, followed by Value/Up, while Gate/Output/Down are the most impactful. This reflects their functional roles: Query/Key support attention scoring, whereas the others play a more direct role in transforming and retaining learned representations.
@@ 4.4 All About Efficiency
We benchmark training time across multiple settings in Table 9 (Appendix G). ABBA has comparable training time to LoRA, with only a $\approx 2-3\%$ overhead, primarily due to the additional computation introduced by the Hadamard product. We clarify that the initialization itself is highly efficient, as we compute the truncated SVD using torch.svd_lowrank. This computation takes less than one second for the entire model, even for the largest LLMs used in our experiments.
@@ 5 Conclusion
6 Limitations
Although we do not evaluate ABBA on ViTs or VLMs, our method is architecture-agnostic and should extend well to such models; we leave this to future work. While ABBA adapters can be merged into the base model to enable zero-overhead inference, serving many adapters concurrently may introduce additional compute due to the Hadamard structure. Lastly, we do not investigate the effect of merging multiple ABBA adapters within a single model. In LoRA, such composition can lead to destructive interference; whether similar effects arise with ABBA remains an open question.
@@ Appendix A Related Work
PEFT methods adapt large pretrained models to downstream tasks by training a small number of additional parameters while keeping the base model frozen. Among these, LoRA [10] is widely adopted for its simplicity and effectiveness, modeling the update $\Delta W$ as a low-rank product $BA$, thereby reducing trainable parameters significantly. Several extensions enhance LoRA along different axes. QLoRA [38] and QA-LoRA [39] combines quantization with LoRA to reduce memory footprint; AdaLoRA [40] allocates rank budgets dynamically across layers. LoRA-XS [12] inserts a small trainable core between frozen adapters to improve compression, while LoRA-Pro [14] and LoRA-SB [15] optimize adapters to better approximate full fine-tuning gradients. Other variants modify structure or initialization. VeRA [41] reuses frozen adapters across layers with task-specific scaling vectors; DoRA [11] applies low-rank updates only to the direction of pretrained weights. PiSSA [26] initializes adapters using top singular vectors, while rsLoRA [19] proposes scale-aware initialization to improve stability. Despite all these differences, these methods share a core principle: they represent $\Delta W$ using a low-rank structure, enabling adaptation under tight compute and memory constraints. LoRA-based methods have also been applied in other domains, such as federated fine-tuning [42, 21, 43, 44].
@@
While low-rank PEFT methods offer strong efficiency–performance tradeoffs, they can underperform in tasks requiring high-rank updates. This limitation has driven recent efforts to move beyond purely low-rank parameterizations. HiRA [13] achieves high-rank updates by taking the Hadamard (elementwise) product of the pretrained matrix $W_{0}$ with a low-rank adapter $BA$, leveraging the fact that such a product can increase effective rank without increasing parameter count. MoRA [45] instead learns a full-rank update through input compression and activation decompression, while KronA [46] uses Kronecker products between adapters to boost representational capacity. ReLoRA [47] uses multiple low-rank updates to approximate a final higher-rank update. Other approaches explore elementwise structure for different purposes: FLoRA [42] modulates intermediate activations per task using Hadamard products, and PACE [48] injects multiplicative noise during adaptation. These trends reflect a broader shift toward structured, high-rank updates for improved expressivity.
@@
Appendix B Why Does No Closed-Form Reconstruction Solution Exist for ABBA?
@@
to decouple along singular directions, as guaranteed by the Eckart–Young–Mirsky theorem [16, 17].
@@ Appendix C Why Does LoRA-Style Initialization of Both Adapter Pairs Fail?
Notice that $\frac{\partial\left(B_{1,il}A_{1,lj}\right)}{\partial A_{1,pq}}=0$ when $l\neq p$ and $j\neq q$ else $=B_{1,ip}$. This means we can rewrite the above summation as
@@
Again notice that $\frac{\partial\left(B_{1,il}A_{1,lj}\right)}{\partial B_{1,pq}}=0$ when $i\neq p$ and $l\neq q$ else $=A_{1,qj}$. This means we can rewrite the above summation as
@@ D.1 Proof of Theorem 1: Khatri–Rao Factorization
Theorem (Khatri–Rao Factorization [18]).
Let $B_{1}A_{1},B_{2}A_{2}\in\mathbb{R}^{m\times n}$ Then, $(B_{1}A_{1})\odot(B_{2}A_{2})=\underbrace{(B_{1}\odot_{r}B_{2})}_{m\times r_{1}r_{2}}\underbrace{(A_{1}^{\top}\odot_{r}A_{2}^{\top})^{\top}}_{r_{1}r_{2}\times n},$ where $\odot_{r}$ denotes the row-wise Khatri–Rao product.
@@
[…]
@@
D.2 Proof of Theorem 2: Effective Scaling for ABBA Update
Theorem (Effective Scaling for ABBA Updates).
To maintain training stability and match the update magnitude of LoRA, the ABBA update $\Delta W=s_{\text{ABBA}}\cdot(B_{1}A_{1}\odot B_{2}A_{2})$ should use the scaling factor $s_{\text{ABBA}}=\frac{\alpha_{\text{LoRA}}^{2}}{\sqrt{r_{1}r_{2}}},$ where $\alpha_{\text{LoRA}}$ is the standard LoRA scaling coefficient and $r_{1},r_{2}$ are the ranks of the two ABBA adapter pairs.
Proof.
We assume that each entry in the low-rank matrices $B_{i}$ and $A_{i}$ is independently drawn from a zero-mean distribution with variance $\sigma^{2}$. Therefore,
$\mathbb{E}[B_{i,jk}^{2}]=\sigma^{2},\quad\mathbb{E}[A_{i,kj}^{2}]=\sigma^{2}.$
We now compute the variance of a single element of the matrix product $B_{i}A_{i}$. For the first adapter pair:
$\mathbb{E}\left[[B_{1}A_{1}]_{ij}^{2}\right]$ $=\mathbb{E}\left[\left(\sum_{k=1}^{r_{1}}B_{1,ik}A_{1,kj}\right)^{2}\right]$
$=\sum_{k=1}^{r_{1}}\mathbb{E}[B_{1,ik}^{2}]\cdot\mathbb{E}[A_{1,kj}^{2}]$
$=r_{1}\sigma^{4}.$
Similarly, for the second adapter pair:
$\mathbb{E}\left[[B_{2}A_{2}]_{ij}^{2}\right]=r_{2}\sigma^{4}.$
The ABBA update is defined as:
$\Delta W_{ij}=s_{\text{ABBA}}\cdot(B_{1}A_{1}\odot B_{2}A_{2})_{ij},$
and the variance of this update becomes:
$\mathbb{E}[\Delta W_{ij}^{2}]$ $=s_{\text{ABBA}}^{2}\cdot\mathbb{E}\left[(B_{1}A_{1})_{ij}^{2}\right]\cdot\mathbb{E}\left[(B_{2}A_{2})_{ij}^{2}\right]$
$=s_{\text{ABBA}}^{2}\cdot r_{1}r_{2}\cdot\sigma^{8}$
$=\left(s_{\text{ABBA}}\cdot\sqrt{r_{1}r_{2}}\cdot\sigma^{4}\right)^{2}.$
To ensure stable training, we require this variance to match a desired target $\tau^{2}$. Assuming $\sigma=1$, this simplifies to:
$\tau=s_{\text{ABBA}}\cdot\sqrt{r_{1}r_{2}}\quad\Rightarrow\quad s_{\text{ABBA}}=\frac{\tau}{\sqrt{r_{1}r_{2}}}.$
Next, we determine the appropriate value of $\tau$ by aligning the effective scale of the ABBA update with that of two independently scaled LoRA-style adapters. LoRA typically applies a scaling factor of $\frac{\alpha_{\text{LoRA}}}{\sqrt{r}}$ to each low-rank adapter. Thus, composing two such adapters yields:
$\left(\frac{\alpha_{\text{LoRA}}}{\sqrt{r_{1}}}B_{1}A_{1}\right)\odot\left(\frac{\alpha_{\text{LoRA}}}{\sqrt{r_{2}}}B_{2}A_{2}\right)$ $=\frac{\alpha_{\text{LoRA}}^{2}}{\sqrt{r_{1}r_{2}}}(B_{1}A_{1}\odot B_{2}A_{2})$
$=\frac{\tau}{\sqrt{r_{1}r_{2}}}(B_{1}A_{1}\odot B_{2}A_{2}),$
which implies $\tau=\alpha_{\text{LoRA}}^{2}$. Substituting this into the earlier expression for $s_{\text{ABBA}}$, we obtain:
$s_{\text{ABBA}}=\frac{\alpha_{\text{LoRA}}^{2}}{\sqrt{r_{1}r_{2}}}.$
This completes the proof. ∎
D.3 Proof of Theorem 3: Probabilistic Advantage of ABBA over LoRA
Definition 2 (Hadamard Product Manifold).
We define the Hadamard product manifold $\mathcal{M}_{\mathrm{Had}}\subset\mathbb{R}^{m\times n}$ as the set of all such matrices: $\mathcal{M}_{\mathrm{Had}}:=\left\{(B_{1}A_{1})\odot(B_{2}A_{2})\right\}\equiv\left\{(B_{1}\odot_{r}B_{2})(A_{1}\odot_{r}A_{2})\right\}.$
Theorem (Probabilistic Advantage of ABBA over LoRA).
Let $M=M_{\mathrm{true}}+\sigma Z$, where $M_{\mathrm{true}}\in\mathcal{M}_{\mathrm{Had}}$ lies on the Hadamard-product manifold (space in which ABBA updates lie), and $Z\in\mathbb{R}^{m\times n}$ has entries i.i.d with each distributed $\mathcal{N}(0,1)$. Let $\epsilon_{\mathrm{bias}}^{2}:=\|M_{\mathrm{true}}-M_{\mathrm{SVD},r}\|_{F}^{2}.$ and $\delta=\exp\left(-\frac{\epsilon_{\mathrm{bias}}^{2}}{8\sigma^{2}}\right)$, then, $\mathbb{P}\left(\|M-M_{\mathrm{ABBA},r}\|_{F}^{2}<\|M-M_{\mathrm{SVD},r}\|_{F}^{2}\right)\geq 1-\delta$ (13)
Proof.
Since $M_{\mathrm{true}}$ lies on the Hadamard manifold, we have $M_{\mathrm{ABBA},r}=M_{\mathrm{true}}$. Therefore, the reconstruction error using our ABBA formulation can be written as:
$\|M-M_{\mathrm{ABBA},r}\|_{F}^{2}=\|M-M_{\mathrm{true}}\|_{F}^{2}=\|\sigma Z\|_{F}^{2}=\sigma^{2}\|Z\|_{F}^{2}.$
Similarly, the reconstruction error for the SVD-based approximation is:
$\|M-M_{\mathrm{SVD},r}\|_{F}^{2}$ $=\|M_{\mathrm{true}}+\sigma Z-M_{\mathrm{SVD},r}\|_{F}^{2}$
$=\|M_{\mathrm{true}}-M_{\mathrm{SVD},r}\|_{F}^{2}+\sigma^{2}\|Z\|_{F}^{2}+2\sigma\langle Z,M_{\mathrm{true}}-M_{\mathrm{SVD},r}\rangle$
$=\epsilon_{\mathrm{bias}}^{2}+\sigma^{2}\|Z\|_{F}^{2}+2\sigma\langle Z,M_{\mathrm{true}}-M_{\mathrm{SVD},r}\rangle,$
where we define $\epsilon_{\mathrm{bias}}^{2}:=\|M_{\mathrm{true}}-M_{\mathrm{SVD},r}\|_{F}^{2}$.
We now compute the difference between the two reconstruction errors:
$\Delta$ $:=\|M-M_{\mathrm{SVD},r}\|_{F}^{2}-\|M-M_{\mathrm{ABBA},r}\|_{F}^{2}$
$=\epsilon_{\mathrm{bias}}^{2}+2\sigma\langle Z,M_{\mathrm{true}}-M_{\mathrm{SVD},r}\rangle.$
Since the entries of $Z$ are i.i.d. and distributed as $\mathcal{N}(0,1)$, the inner product $\langle Z,M_{\mathrm{true}}-M_{\mathrm{SVD},r}\rangle$ is normally distributed with mean 00 and variance $\|M_{\mathrm{true}}-M_{\mathrm{SVD},r}\|_{F}^{2}=\epsilon_{\mathrm{bias}}^{2}$. Hence, $\Delta\sim\mathcal{N}\left(\epsilon_{\mathrm{bias}}^{2},4\sigma^{2}\epsilon_{\mathrm{bias}}^{2}\right)$.
We are interested in the probability that ABBA achieves a lower reconstruction error than SVD, i.e., $\mathbb{P}[\Delta>0]$. Applying the Chernoff tail bound:
$\mathbb{P}[\Delta\leq 0]=\mathbb{P}\left[\langle Z,M_{\mathrm{true}}-M_{\mathrm{SVD},r}\rangle<-\frac{\epsilon_{\mathrm{bias}}^{2}}{2\sigma}\right]\leq\exp\left(-\frac{1}{2}\cdot\left(\frac{\epsilon_{\mathrm{bias}}^{2}}{2\sigma\cdot\sqrt{\epsilon_{\mathrm{bias}}^{2}}}\right)^{2}\right)$
$=\exp\left(-\frac{\epsilon_{\mathrm{bias}}^{2}}{8\sigma^{2}}\right)$
$\implies\mathbb{P}[\Delta>0]\geq 1-\delta$
thus completing the proof. ∎
D.4 Proof of Theorem 4: Rank-Stability of ABBA Initialization
Theorem (Rank-Stability of ABBA Initialization).
Let $s_{\text{ABBA}}$ be the scaling factor from Theorem 2, and let the initialization follow Section 2.3. Then the ABBA update satisfies the forward and backward stability conditions of Definition 1, and is thus rank-stable.
Proof.
We follow a similar analysis to that presented in LoRA-GA [49].

### v3

@@ 3.2 Arithmetic Reasoning
We fine-tune Mistral-7B (24) and Gemma-2 9B (25) on a 20K-sample subset of MetaMathQA (35), and evaluate their performance on GSM8K (36) and MATH (37). We insert LoRA adapters into all attention projections (query, key, value, and output) as well as both feedforward layers. We report results in Table 2. ABBA achieves superior performance over all other PEFT approaches across both models, and often outperforms full FT. We hypothesize that this effect arises because ABBA’s structured parameterization implicitly regularizes training and enables more efficient task-specific adaptation than unconstrained full FT.
@@
Method Mistral‑7B Gemma‑2 9B
# Params GSM8K ($\uparrow$) MATH $(\uparrow$) # Params GSM8K ($\uparrow$) MATH $(\uparrow$)
@@ 4.1 Initialization Strategies for ABBA
Initialization of the adapter matrices $B_{1},A_{1}\text{ and }B_{2},A_{2}$ is crucial to ABBA’s performance. A naive LoRA-style initialization, where $B_{1},B_{2}$ are set to zero and $A_{1},A_{2}$ use Kaiming uniform, leads to training failure due to zeroed-out gradients (see Appendix C). To address this, we explore several initialization strategies that combine truncated SVD-based approximations with standard schemes, summarized in Table 3.
Inspired by PiSSA-LoRA (26), one approach is to approximate the base weight $W_{0}$ at initialization. This can be done by initializing one adapter pair using the truncated SVD of $W_{0}$, and the other with scaled constant values (e.g., ones) to prevent gradient explosion (1). Another strategy initializes both adapter pairs using the top-$r/2$ components from the truncated SVD of $\sqrt{W_{0}}$ (2). A variation of this approach assigns the top-$r/2$ components to one adapter pair and the next-$r/2$ to the other, introducing greater representational diversity and yielding slightly improved results (3, 4). We also consider a hybrid strategy, where one adapter pair approximates $W_{0}$ via truncated SVD, and the other follows LoRA-style initialization (Kaiming for $A$, zeros for $B$). This configuration performs best and closely resembles the initialization used in HiRA (5).
Our final method adopts this approach: $B_{1},A_{1}$ are initialized using the truncated SVD of $W_{0}$, while $B_{2},A_{2}$ follow LoRA-style initialization (Ours).
@@
(1) $B_{1},A_{1}\leftarrow\text{top-}r/2\text{ from Trunc. SVD}(W_{0})$, $B_{2},A_{2}\leftarrow\text{Ones (scaled)}$ 57.39 10.88
(2) Both adapters: top-$r/2$ from $\text{Trunc. SVD}(\sqrt{W_{0}})$ 64.53 16.57
(3) First adapter: top-$r/2$, second: next-$r/2$ from $\text{Trunc. SVD}(\sqrt{W_{0}})$ 64.86 17.05
(4) Second adapter: top-$r/2$, first: next-$r/2$ from $\text{Trunc. SVD}(\sqrt{W_{0}})$ 64.79 17.16
(5) $B_{2},A_{2}\leftarrow\text{top-}r/2\text{ from Trunc. SVD}(W_{0})$, $B_{1},A_{1}\leftarrow$ LoRA Init (Zeros, Kaiming) 66.19 18.06
@@ 4.2 Choosing Important Hyperparameters
To empirically validate this, we sweep over a range of $\alpha$ values for Llama-3.2 3B, in Table 4. Consistent with prior PEFT literature, ABBA achieves optimal performance within the typical LoRA scaling range of 163216-32. Additional evidence is provided in Table 7 (Appendix E).
Table 4: Performance comparison across different $\alpha$ values for Llama-3.2 3B.
$\alpha$ Accuracy ($\uparrow$)
@@ 4.3 Placement of ABBA in Transformers
Figure 3 examines the effect of fine-tuning individual transformer components in ABBA, namely Query, Key, Value, Output, Up, Gate, and Down projections. The results show the following: Query/Key contribute the least, followed by Value/Up, while Gate/Output/Down are the most impactful. This reflects their functional roles: Query/Key support attention scoring, whereas the others play a more direct role in transforming and retaining learned representations.
@@ 4.4 All About Efficiency
We benchmark training time across multiple settings in Table 9 (Appendix G). ABBA has comparable training time to LoRA, with only a $\approx 2-3\%$ overhead, primarily due to the additional computation introduced by the Hadamard product. We clarify that the initialization itself is highly efficient, as we compute the truncated SVD using torch.svd_lowrank. This initialization step takes less than one second for the entire model, even for the largest LLMs used in our experiments.
@@ Appendix A Related Work
PEFT methods adapt large pretrained models to downstream tasks by training a small number of additional parameters while keeping the base model frozen. Among these, LoRA (10) is widely adopted for its simplicity and effectiveness, modeling the update $\Delta W$ as a low-rank product $BA$, thereby reducing trainable parameters significantly. Several extensions enhance LoRA along different axes. QLoRA (38) and QA-LoRA (39) combines quantization with LoRA to reduce memory footprint; AdaLoRA (40) allocates rank budgets dynamically across layers. LoRA-XS (12) inserts a small trainable core between frozen adapters to improve compression, while LoRA-Pro (14) and LoRA-SB (15) optimize adapters to better approximate full fine-tuning gradients. Other variants modify structure or initialization. VeRA (41) reuses frozen adapters across layers with task-specific scaling vectors; DoRA (11) applies low-rank updates only to the direction of pretrained weights. PiSSA (26) initializes adapters using top singular vectors, while rsLoRA (21) proposes scale-aware initialization to improve stability. Despite all these differences, these methods share a core principle: they represent $\Delta W$ using a low-rank structure, enabling adaptation under tight compute and memory constraints. LoRA-based methods have also been applied in other domains, such as federated fine-tuning (42, 20, 43, 44).
@@
While low-rank PEFT methods offer strong efficiency–performance tradeoffs, they can underperform in tasks requiring high-rank updates. This limitation has driven recent efforts to move beyond purely low-rank parameterizations. HiRA (13) achieves high-rank updates by taking the Hadamard (elementwise) product of the pretrained matrix $W_{0}$ with a low-rank adapter $BA$, leveraging the fact that such a product can increase effective rank without increasing parameter count. MoRA (45) instead learns a full-rank update through input compression and activation decompression, while KronA (46) uses Kronecker products between adapters to boost representational capacity. ReLoRA (47) uses multiple low-rank updates to approximate a final higher-rank update. Other approaches explore elementwise structure for different purposes: FLoRA (42) modulates intermediate activations per task using Hadamard products, and PACE (48) injects multiplicative noise during adaptation. These trends reflect a broader shift toward structured, high-rank updates for improved expressivity.
@@
Appendix B Why Does No Closed‑Form Reconstruction Solution Exist for ABBA?
@@
to decouple along singular directions, as guaranteed by the Eckart–Young–Mirsky theorem (16, 17).
@@ Appendix C Why Does LoRA-Style Initialization of Both Adapter Pairs Fail?
Notice that $\frac{\partial\left(B_{1,il}A_{1,lj}\right)}{\partial A_{1,pq}}=0$ when $l\neq p$ and $j\neq q$ else $B_{1,ip}$. This means we can rewrite the above summation as
@@
Again notice that $\frac{\partial\left(B_{1,il}A_{1,lj}\right)}{\partial B_{1,pq}}=0$ when $i\neq p$ and $l\neq q$ else $A_{1,qj}$. This means we can rewrite the above summation as
@@ D.1 Proof of Theorem 1: Khatri–Rao Factorization
Theorem (Khatri–Rao Factorization (18)).
Let $B_{1}A_{1},B_{2}A_{2}\in\mathbb{R}^{m\times n}$ . Then, $(B_{1}A_{1})\odot(B_{2}A_{2})=\underbrace{(B_{1}\odot_{r}B_{2})}_{m\times r_{1}r_{2}}\underbrace{(A_{1}^{\top}\odot_{r}A_{2}^{\top})^{\top}}_{r_{1}r_{2}\times n},$ where $\odot_{r}$ denotes the row-wise Khatri–Rao product.
@@
$[U\odot_{r}V]_{1}$ $:=[1\cdot V_{1},\;2\cdot V_{1}]=\begin{bmatrix}3&4&6&8\end{bmatrix}$,
$[U\odot_{r}V]_{2}$ $:=[3\cdot V_{2},\;4\cdot V_{2}]=\begin{bmatrix}15&18&20&24\end{bmatrix}$.
@@
D.2 Proof of Theorem 2: Rank-Stability of ABBA
Theorem (Rank-Stability of ABBA).
Let the ABBA update be $\Delta W=s_{\text{ABBA}}\,(B_{1}A_{1})\odot(B_{2}A_{2}),$ with $B_{1},A_{1},B_{2},A_{2}$ independent, mean-zero random matrices with finite variance. Then, the update satisfies the stability conditions of Definition 1 if and only if: $s_{\text{ABBA}}\in\Theta\!\left(\tfrac{1}{\sqrt{r_{1}r_{2}}}\right).$